Actions from previous week:

o    Check the availability of a beam pipe to allow the installation of only two CLIC cavity BPM      > ongoing
o    Digital camera 620 needs to be relocated to the right server (Stefano is already working on it).      > completed.
o    Follow-up the compression low/ trigger issues.      > the origin was identified and it's being fixed. To be checked.

o   Check with vacuum team the noise of the plasma lens turbo pump.